OpenStudy (selena2345):

Fill in the missing number 84/6- Missing Number x 3 =11

OpenStudy (anonymous):

84/6-m.3=11 84/6-11=3m 18/6=3m 3=3m 1=m m:missing number

OpenStudy (darkprince14):

is the 1st equation m*3?

OpenStudy (anonymous):

yep

OpenStudy (selena2345):

84/6-Missing Number x 3 =11

OpenStudy (darkprince14):

was three multiplied to m or is it multiplied to the fraction??

OpenStudy (darkprince14):

@Selena2345

OpenStudy (selena2345):

yes

OpenStudy (calculusfunctions):

@Selena2345, I thought you said you understood after the last problem which was exactly the same. Were you being less than honest?

OpenStudy (darkprince14):

if it is multiplied to the fraction, it's like this: (84/6-m) (3) = 11 then, divide 3 from both sides(the same as mutiplying 1/3 to both sides) \[\frac{ 1 }{3 } \left( \frac{ 84 }{ 6-m} \right)\left( 3 \right) = \frac{ 1 }{ 3 } \left( 11 \right)\] so, it'll be like this \[\frac{ 84 }{ 6-m } = \frac{ 11 }{ 3 }\] cross multiply (6-m)(11) = (84)(3) 66 - 11m = 252 -11m = 252 - 66 -11m = 186 \[\frac{ -11m }{ -11 } = \frac{ 186 }{ -11 }\]

OpenStudy (darkprince14):

don't know if it's righ... t try checking it.. \[m = \frac{-186? }{ 11 }\]

OpenStudy (selena2345):

thank you so much your the best

OpenStudy (darkprince14):

@calculusfunctions Is my solution right?? i kinda doubt my answers..

OpenStudy (selena2345):

your answer is very good and smart

OpenStudy (darkprince14):

thanks..

OpenStudy (selena2345):

your welcome

OpenStudy (calculusfunctions):

@darkprince, your first answer of m = 1 is perfect! Your second answer however, is incorrect. The question is\[\frac{ 84 }{ 6 }-3m =11\]Therefore as I said, your first response was perfect!

OpenStudy (darkprince14):

sorry..i got the wrong picture.. if you simplified the fraction, you'll get 14-3m = 11 14 - 11 = 3m 3 =3m 1 = m